### v1

# How Easy is It to Fool Your Multimodal LLMs? An Empirical Analysis on Deceptive Prompts

Abstract
The remarkable advancements in Multimodal Large Language Models (MLLMs) have not rendered them immune to challenges, particularly in the context of handling deceptive information in prompts, thus producing hallucinated responses under such conditions. To quantitatively assess this vulnerability, we present MAD-Bench, [^1] Short for MultimodAl Deception Benchmark. a carefully curated benchmark that contains 850 test samples divided into 6 categories, such as non-existent objects, count of objects, spatial relationship, and visual confusion. We provide a comprehensive analysis of popular MLLMs, ranging from GPT-4V, Gemini-Pro, to open-sourced models, such as LLaVA-1.5 and CogVLM. Empirically, we observe significant performance gaps between GPT-4V and other models; and previous robust instruction-tuned models, such as LRV-Instruction and LLaVA-RLHF, are not effective on this new benchmark. While GPT-4V achieves 75.02% accuracy on MAD-Bench, the accuracy of any other model in our experiments ranges from 5% to 35%. We further propose a remedy that adds an additional paragraph to the deceptive prompts to encourage models to think twice before answering the question. Surprisingly, this simple method can even double the accuracy; however, the absolute numbers are still too low to be satisfactory. We hope MAD-Bench can serve as a valuable benchmark to stimulate further research to enhance models’ resilience against deceptive prompts.
How Easy is It to Fool Your Multimodal LLMs?
An Empirical Analysis on Deceptive Prompts
Yusu Qian, Haotian Zhang, Yinfei Yang, Zhe Gan Apple {yusuqian,haotian.zhang2,yinfeiy,zhe.gan}@apple.com
1 Introduction
Figure 1: How easy is it to fool your multimodal LLMs? Our study found that multimodal LLMs, such as LLaVA-1.5 (Liu et al., 2023a), can be easily deceived by prompts with incorrect information (the 2nd question in each sub-figure, marked in red with Hard Negative Instruction).
Recent advancements in Multimodal Large Language Models (MLLMs) (Liu et al., 2023b, a; Wang et al., 2023c; You et al., 2024; Bai et al., 2023b; Liu et al., 2024; Zhu et al., 2024), exemplified by models like GPT-4V(ision) (OpenAI, 2023) and Gemini (Team, 2023), mark a significant milestone in the evolution of AI, extending the capabilities of large language models to the realm of visual understanding and interaction.
However, the sophistication of MLLMs brings with it unique challenges, notably, hallucination. Current studies (Liu et al., 2024; Lee et al., 2023; Yin et al., 2023) have been actively exploring solutions to mitigate hallucination, especially when the model tries to generate long responses. However, there still remains a notable gap in the literature: no work has yet been conducted to focus on comprehensively studying the robustness of MLLMs when confronted with deceptive information in the prompts. [^2] LRV-Instruction (Liu et al., 2023a) is the pioneering work in this direction, while we aim to provide a more comprehensive evaluation with hard negative instructions. Please see Section 2 for a more detailed discussion on related work. Our work aims to fill in this gap. This issue is particularly critical, as it pertains to the reliability and trustworthiness of these models in real-world applications (Liu et al., 2023c), and holds substantial importance for the ongoing development and deployment of such AI systems.
To this end, we present [Uncaptioned image] MAD-Bench, a carefully curated benchmark that contains 850 image-prompt pairs spanning across six deception categories, to systematically examine how MLLMs resolve the conflicts when facing inconsistencies between text prompts and images. We provide a comprehensive analysis of popular MLLMs, ranging from GPT-4V (OpenAI, 2023), Gemini-Pro (Team, 2023), to open-sourced models, such as LLaVA-1.5 (Liu et al., 2023a) and CogVLM (Wang et al., 2023c). The evaluation is fully automated via the use of GPT-4. Results shed light on how vulnerable MLLMs are in handling deceptive instructions. For example, Figure 1 illustrates how sensitive LLaVA-1.5 (Liu et al., 2023a) is to the factualness of the input prompt and its consistency with the image. When asked “is there a cat in the image?”, LLaVA-1.5 can successfully identify there is no cat; but when prompted with “what color is the cat in the image?”, the model will imagine there is a cat inside. Empirically, we observe that GPT-4V suffers much less when compared with all the other MLLMs; however, the performance is still not ideal (GPT-4V vs. others: 75% vs. 5%-35% accuracy). Further, previous models that aim to mitigate hallucinations, such as LRV-Instruction (Liu et al., 2024) and LLaVA-RLHF (Sun et al., 2023b), are not effective on this new benchmark.
Finally, we provide a simple remedy to boost performance, which was surprisingly found to be effective to double the models’ accuracy. Specifically, we carefully design a system prompt in the form of a long paragraph to be prepended to the existing prompt, to encourage the model to think carefully before answering the question. This simple approach boosts the accuracy of LLaVA-1.5 from 10.42% to 20.56% (similar boosts for other models); however, the absolute numbers are still too low to be satisfactory. Further research is needed to study how to match GPT-4V’s performance (75.02%).
Our contributions are summarized as follows. ($i$) We construct MAD-Bench, a new benchmark to comprehensively evaluate MLLMs on their capability to resist deceiving information in the prompt. ($ii$) We provide a detailed analysis of popular MLLMs, and list some common causes for incorrect responses. ($iii$) We provide a simple remedy to boost performance via the careful design of a system prompt. MAD-Bench will be open-sourced, and we hope this benchmark can serve as a useful resource to stimulate further research to enhance models’ resilience against deceptive prompts.
2 Related Work
Multimodal Large Language Models (MLLMs).
MLLM has become an increasingly hot research topic. Early models primarily focused on large-scale image-text pre-training (Wang et al., 2022b, a; Chen et al., 2022, 2023c; Li et al., 2023c; Driess et al., 2023; Huang et al., 2023; Awadalla et al., 2023; Laurençon et al., 2023). Among them, Flamingo (Alayrac et al., 2022) pioneered the integration of a CLIP image encoder with LLMs through gated cross-attention blocks, showcasing emergent multimodal in-context few-shot learning capabilities, via pre-training over millions of image-text pairs and interleaved image-text datasets (Zhu et al., 2023).
On the other hand, recent research has focused on visual instruction tuning (Zhu et al., 2024; Li et al., 2023a; Ye et al., 2023a; Li et al., 2023b; Chen et al., 2023b). Prominent examples include LLaVA(-1.5) (Liu et al., 2023b, a), InstructBLIP (Dai et al., 2023), Qwen-VL (Bai et al., 2023a), CogVLM (Wang et al., 2023c), Emu2 (Sun et al., 2023a), SPHINX (Lin et al., 2023), to name a few. Besides text response generation, recent works have also enabled MLLMs for referring and grounding (Peng et al., 2023a; Chen et al., 2023a; You et al., 2024; Wang et al., 2023d), image segmentation (Lai et al., 2023; Zhang et al., 2023), image editing (Fu et al., 2023b), image generation (Koh et al., 2023; Sun et al., 2023a), etc.
The release of proprietary systems like GPT-4V (OpenAI, 2023) and Gemini (Team, 2023) has elevated the research of MLLMs to new heights. Since GPT-4V’s release, researchers have been exploring its capabilities as well as weaknesses (Zhou et al., 2023; Li et al., 2023f; Liu et al., 2023e; Yang et al., 2023; Cui et al., 2023). As MLLMs become stronger, the development of more challenging benchmarks is essential to push the boundaries of what these models can achieve. In this work, we aim to design a new benchmark to evaluate MLLMs’ resilience against deceptive prompts.
Figure 2: Examples of deceptive prompts used in the proposed MAD-Bench with example model responses.
Hallucination in MLLMs.
Below, we first discuss hallucination in LLMs, and then focus on hallucination in MLLMs.
Existing work on mitigating hallucination in LLMs can be roughly divided into two categories: ($i$) prompt engineering (Si et al., 2023; Cheng et al., 2023; Ji et al., 2023; Jones et al., 2023; Mündler et al., 2023; Vu et al., 2023), and ($ii$) model enhancement (Li et al., 2023d; Chuang et al., 2023; Shi et al., 2023; Elaraby et al., 2023; Tian et al., 2024; Qiu et al., 2023; Leng et al., 2023). These studies laid solid foundations for understanding the causes of hallucinations, such as over-reliance on context, or training data biases.
Similarly, hallucination in MLLMs is also growing to be an important research topic (Liu et al., 2024). There are various categories of hallucinations, such as describing objects that are non-existent in the input image, misunderstanding the spatial relationship between objects in the image, and counting objects incorrectly (Liu et al., 2023d). The two main causes of hallucination in MLLMs found in existing work apart from the potential issues with training data include ($i$) limitations in correctly understanding input images, and ($ii$) language model bias (Wang et al., 2023b). Various methods have been proposed to mitigate hallucination in MLLMs (Lee et al., 2023; Yin et al., 2023; Sun et al., 2023b; Wang et al., 2023a; Liu et al., 2024; Zhai et al., 2023; Zhou et al., 2024; Gunjal et al., 2024; Liu et al., 2023b).
Furthermore, various benchmarks have been proposed to evaluate hallucination in MLLMs. Specifically, POPE (Li et al., 2023e), M-HalDetect (Gunjal et al., 2024), and GAVIE (Liu et al., 2024) evaluated object hallucination. HallusionBench (Guan et al., 2023) evaluated both visual and language hallucination. MMHal-Bench (Sun et al., 2023b) evaluated hallucination in more aspects including relations, attributes, environments, etc. Bingo (Cui et al., 2023) studied hallucination in terms of bias and interference in GPT-4V (OpenAI, 2023).
In this work, we aim to study how easy it is to use deceptive prompts that contain information inconsistent with the image to mislead MLLMs to generate responses with hallucination. Note, that we are not the first to study this. A similar model behavior is called “sycophancy” in the LLM literature (Sharma et al., 2023). Fu et al. (2023a) and Liu et al. (2023a) also constructed prompts with deceiving information to test model robustness. Deceptive prompts are termed “negative instructions” in LRV-Instruction (Liu et al., 2023a) and “text-to-image interference” in the Bingo benchmark (Cui et al., 2023). Different from them, we comprehensively study MLLMs’ ability to handle deceptive prompts in multiple categories. Unlike previous studies (Fu et al., 2023a; Liu et al., 2023a) which primarily used “Is/Are/Can” questions, we found that it is relatively easy for state-of-the-art MLLMs to counter deceptive information in such formats. Consequently, we shifted our focus to questions beginning with “What”, “How”, “Where”, etc., to provide a more challenging and insightful evaluation.
3 MAD-Bench
In this section, we present MAD-Bench, introduce how we collect [^3] The authors are responsible for collecting the data, Apple Inc is not. deceptive image-prompt pairs, as well as our evaluation method.
3.1 Deception Categories
MAD-Bench encompasses six distinct categories of 850 image-prompt pairs designed to test the resilience of MLLMs against deceptive prompts. Table 1 provides the statistics of each category, and Figure 2 shows examples of deceptive prompts. The selected categories are partly inspired by Liu et al. (2023d). Below, we detail each category.
Figure 3: Examples of image-prompt pairs in the Visual Confusion category of MAD-Bench.
Count of Object. This category intentionally cites an incorrect quantity of visible objects in the image. A response fails this test if it asserts the presence of $m$ instances of an object ‘A’ when, in reality, a different number $n$ of object ‘A’ is present — $n$ being distinct from $m$ and not zero. The images for this and the subsequent four categories are sourced from COCO 2017 (Lin et al., 2015). Using a public dataset sometimes brings concerns about data leakage. In our case, given the special nature of our deceptive prompts to be introduced in the next section, this will not be a problem. An accurate response would either challenge the prompt’s inconsistency with the visual data and abstain from speculating on absent information, or seek further clarification to resolve any uncertainties.
Deception Category Count Image Source
Count of Object 188 COCO 2017
Non-existent Object 244 COCO 2017
Object Attribute 136 COCO 2017
Scene Understanding 122 COCO 2017
Spatial Relationship 132 COCO 2017
Visual Confusion 28 In the Wild
Table 1: Statistics of the 850 image-prompt pairs in MAD-Bench.
Non-existent Object. Here, the prompts query about objects absent from the image. Failure occurs when a response acknowledges these non-existent objects as present.
Object Attribute. This category includes prompts that inaccurately describe visible objects’ attributes. A response fails if it attributes these incorrect characteristics to the actual objects in the image.
Figure 4: Illustration of the process of generating deceptive prompts in the non-existent object category using GPT-4 and COCO ground-truth captions.
Scene Understanding. This category involves prompts that inaccurately describe the scene encapsulating the objects in the image. A response that falls into error here can be one that accurately identifies the actions of the objects but misconstrues the scene or setting in alignment with the deceptive prompt.
Spatial Relationship. This category presents prompts that incorrectly specify the spatial dynamics between objects that do indeed exist within the image. A misstep in this category arises when a response correctly recognizes the objects but misrepresents their spatial relations.
Visual Confusion. This category is different from the previous ones by employing both the prompts and the images as instruments of deception, often deceptive even to the human eye. This category includes three types of images: ($i$) those depicting naked-eye 3D paintings or screens, ($ii$) visual dislocation photography, and ($iii$) mirror reflections. Figure 3 shows an example image-prompt pair  [^4] Photo credit to Braga last1 and Tiago Silva. in each category. Here, the prompts paired with the 3D paintings or screens aim to deceive the MLLMs by portraying the objects in the two-dimensional artwork as three-dimensional entities. With visual dislocation photography, the prompts reinforce the optical illusions present in the images. Lastly, the prompts associated with mirror reflections attempt to deceive the MLLMs into interpreting reflections as tangible objects.
3.2 Prompt Generation Method
The process of creating deceptive prompts was automated by employing GPT-4, leveraging the ground-truth captions from the COCO dataset (Lin et al., 2015). We chose not to use GPT-4V for this task, as we later also evaluated GPT-4V on this benchmark, and empirically, employing GPT-4 is already enough for this task. To guide GPT-4 in generating questions that would intentionally mislead MLLMs within the specified categories, we crafted tailored prompts. These guiding prompts are provided in Appendix A.2, from Figure 18 to 20. The process is illustrated in Figure 4, using an example in the non-existent object category. Bounding box information is not used as part of the prompt sent to GPT-4, as empirically, we observed that it does not contribute to further improving the quality of generated prompts in our deceptive categories. Following the generation of these deceptive questions, a rigorous manual filtering process is followed to ensure that each question adheres to its category’s deceptive criteria and maintains relevance to its associated image.
3.3 Response Evaluation Method
We use GPT-4 to evaluate generated responses from 10 models, including ($i$) 6 open-sourced models: LLaVA-1.5 (Liu et al., 2023a), InstructBLIP (Dai et al., 2023), Ferret (You et al., 2024), Kosmos-2 (Peng et al., 2023b), mPLUG-Owl2 (Ye et al., 2023b), and CogVLM (Wang et al., 2023c), ($ii$) 2 additional open-sourced models that aim to reduce hallucination: LLaVA-RLHF (Sun et al., 2023b) and LRV-V1 (Liu et al., 2024), and ($iii$) 2 state-of-the-art proprietary systems: Gemini-Pro (Team, 2023) and GPT-4V (OpenAI, 2023).
The number of images in the Visual Confusion category is relatively small, while most of them contain humans, so we did not evaluate Gemini in this category as it cannot generate responses for images containing humans. The effect of this on other categories is neglectable. Mirroring the prompt generation method, we design specific prompts for each deceptive category to critically assess the responses. Our primary metric of evaluation is binary, focused strictly on whether the response has been misled, without considering other qualitative aspects such as helpfulness. These prompts for model evaluation are provided in Appendix A.3.
Figure 5: Example failure cases of GPT-4V (OpenAI, 2023).
To verify the accuracy of GPT-4’s automated evaluation, we randomly select 500 responses spanning the various models and deceptive categories for a manual accuracy check. This validation process yielded a 97.0% concordance rate with the outcomes of human evaluation, underlining the reliability of our approach.
4 Experiments
Models
Count of
Object
Non-existent
Object
Object
Attribute
Scene
Understanding
Spatial
Relationship
Visual
Confusion
Meta
Average
M1 Ferret (You et al., 2024) 10.16% 4.94% 5.93% 9.92% 2.29% 7.14% 6.63%
InstructBLIP (Dai et al., 2023) 0.53% 9.47% 11.11% 7.43% 3.05% 21.43% 6.86%
Kosmos-2 (Peng et al., 2023b) 5.34% 0.41% 21.48% 16.53% 3.05% 3.57% 7.70%
LLaVA-1.5 (Liu et al., 2023b) 4.81% 12.35% 11.11% 25.62% 1.53% 3.57% 10.42%
mPLUG-Owl2 (Ye et al., 2023b) 8.02% 22.22% 18.52% 38.84% 9.16% 3.58% 18.23%
CogVLM (Wang et al., 2023c) 14.97% 52.67% 34.07% 33.88% 18.32% 21.43% 32.30%
M2 LRV-V1 (Liu et al., 2024) 5.88% 7.00% 17.78% 43.80% 7.63% 21.43% 14.33%
LLaVA-RLHF (Sun et al., 2023b) 9.63% 14.00% 12.59% 38.02% 3.82% 28.57% 15.15%
M3 Gemini-Pro (Team, 2023) 13.37% 20.99% 38.52% 25.62% 14.50% ${}^{\dagger}$ 21.79%
GPT-4V (OpenAI, 2023) 71.66% 81.07% 71.11% 94.21% 50.38% 96.43% 75.02%
Table 2: Main results on MAD-Bench. M1 denotes open-sourced models. M2 denotes additional open-sourced models that aim to reduce hallucination. M3 denotes state-of-the-art proprietary systems. ($\dagger$) Gemini-Pro cannot respond to images containing humans, and most images in the Visual Confusion category contain humans, thus we skip the evaluation of Gemini-Pro on this category. No response due to humans in the image in the other five categories only occurred six times, and we neglected those when evaluating Gemini’s accuracy. The meta average of accuracy is weighted by the amount of data in each category.
4.1 Main Results
Results are summarized in Table 2. Notably, GPT-4V’s accuracy in the Scene Understanding and Visual Confusion categories is remarkably higher than the others, with over 90% accuracy. This indicates a substantial advancement in GPT-4V’s ability to resist deceptive information. Even LRV-V1 (Liu et al., 2024), whose training data includes negative instructions specifically designed to reduce hallucination in model responses, does not have satisfactory performance in face of deceptive information in our prompts. This is likely because ($i$) the way we design our prompts presents a larger challenge to MLLMs than the “Is/Are/Can”-style negative instructions in Liu et al. (2024), as our prompts are designed intentionally to sound confident in the deceptive information, and ($ii$) their method doesn’t sufficiently generate diverse enough negative prompts.
Interestingly, we observe that models that support bounding box input and output (i.e., Ferret (You et al., 2024) and Kosmos-2 (Peng et al., 2023b)) achieve poor performance on this benchmark. We hypothesize that these models attempt to ground objects as best as they can as they are trained on positive data, therefore, they tend to ground non-existent objects as they are mentioned in the prompts, thus performing poorer than other models on our benchmark. Example responses from each model are provided in Appendix A.1 from Figure 9-15.
Figure 6: Examples of mistakes made by Ferret (You et al., 2024) in face of deceptive prompts. We use Ferret responses for these examples here, as Ferret provides bounding boxes that unveil error types straightforwardly.
Overall, GPT-4V demonstrates superior performance across all metrics compared to the other models. GPT-4V has a more sophisticated understanding of visual data and is less prone to being misled by inaccurate information. This could be attributed to more advanced training, better architecture, or more sophisticated data processing capabilities. The results underscore the potential of GPT-4V in applications where accuracy in interpreting visual and contextual data is critical, despite the challenges of deceptive information. That being said, GPT-4V still fails in many cases, with two examples shown in Figure 5.
4.2 Detailed Analysis
Our examination of how the model reacts to deceptive prompts has uncovered a range of common causes for incorrect responses. Figure 6 illustrates representative instances of errors corresponding to each identified category of mistakes, using Ferret (You et al., 2024) as the running example.
Inaccurate object detection. State-of-the-art MLLMs generally perform well in object detection if not fed deceptive prompts. However, in face of a deceptive prompt mentioning objects invisible in the image, these models may erroneously identify other objects as those mentioned in the prompt.
Redundant object identification. A notable issue arises when the model fails to accurately discern distinct objects referenced in the prompt within the image. This often results in the erroneous identification of a single object as multiple entities, leading to repetitive descriptions as if there were several distinct objects present.
Inference of non-visible objects. The model occasionally attributes characteristics or actions to objects that are not visible in the image. This phenomenon appears to stem from the language model’s reliance on its internal knowledge base to fabricate descriptions for objects mentioned in the prompt but absent in the visual data. Intriguingly, this occurs even when the model does not question the accuracy of its visual recognition capabilities, confidently affirming its findings while simultaneously describing non-existent objects.
Inconsistent reasoning. Throughout the response generation process, we observe the MLLMs oscillating between adhering to the deceptive information in the prompts and relying on their recognition of the actual content in the input image. Sentences in the generated response contradict each other. This inconsistency highlights a fundamental challenge in the model’s decision-making process.
5 A Simple Remedy to Boost Performance
In this section, we introduce a simple yet effective method to enhance the robustness of MLLMs against deceptive prompts while ensuring output alignment with the corresponding input images. This enhancement is realized through the integration of an additional paragraph into the system’s prompt, which is either prepended directly to the existing prompt, or incorporated differently, depending on the specific model.
Figure 7: The additional paragraph prepended to the deceptive prompts to boost performance.
Figure 8: Model responses of mPLUG-Owl2 (Ye et al., 2023b), Gemini-Pro (Team, 2023), and LLaVA-1.5 (Liu et al., 2023b) before and after modifying the test prompt. We add the (*) symbol to the original model name to denote the enhanced model.
Models
Count of
Object
Non-existent
Object
Object
Attribute
Scene
Understanding
Spatial
Relationship
Visual
Confusion
Meta
Average
LLaVA-1.5* 6.38% (+1.57%) 24.69% (+12.34%) 32.59% (21.48%) 24.79% (-0.83%) 17.56% (16.03%) 17.86% (14.29%) 20.56% (+10.14%)
LLaVA-RLHF* 8.56% (-1.07%) 33.61% (+19.61%) 26.67% (+14.08%) 22.13% (-15.89%) 19.08% (+15.26%) 32.14% (+3.57%) 23.01% (+7.86%)
mPLUG-Owl2* 20.32% (+12.30%) 76.54% (+54.32%) 46.67% (+24.15%) 60.33% (+21.49%) 26.72% (+17.56%) 42.86% (+39.28%) 48.15% (+29.92%)
Gemini-Pro* 31.55% (+18.18%) 65.43% (+44.44%) 46.67% (+8.15%) 58.68% (+33.06%) 36.64% (+22.14%) ${}^{\dagger}$ 48.95% (+27.16%)
GPT-4V* 82.35% (+10.69%) 82.72% (+1.65%) 88.89% (+17.78%) 95.90% (+1.69%) 75.57% (+25.19%) 92.86% (-3.57%) 84.74% (+9.72%)
Table 3: Results on MAD-Bench after modifying the test prompt. ($\dagger$) Gemini-Pro cannot respond to images containing humans, and most images in the Visual Confusion category contain humans, thus we skip the evaluation of Gemini-Pro in this category. This simple approach is only tested on models that support and suit this method. The numbers outside of the brackets denote the absolute accuracy, and the numbers inside the brackets denote the performance gain compared to the original models.
We composed this additional paragraph with the help of GPT-4, as shown in Figure 7. It encourages the model to think twice or step by step before answering the question. The model performance after the incorporation of this prompt modification is presented in Table 3. For example, for LLaVA-1.5, it boosts the performance by +10.14%, though the absolute accuracy is still too low to be satisfactory. For GPT-4V, which already achieves an accuracy of 75.02%, using the proposed simple method can further boost the accuracy to 84.74%. Figure 8 provides examples to illustrate the capability of mPLUG-Owl2 (Ye et al., 2023b), LLaVA-1.5 (Liu et al., 2023b) and Gemini-Pro (Team, 2023) to withstand deceptive prompts when supported by modifications made to the test prompt.
Overall, the addition of prompts to resist deceptive information appears to bolster the performance, enabling MLLMs to handle deception better and interpret scenes more accurately. This suggests that strategic prompt design could be a valuable approach to enhancing the robustness of AI models against attempts to mislead or confuse them. Note, that the implementation has not been fully optimized, and some MLLMs do not support this method due to reasons such as limitation of input sequence length. The goal here is to demonstrate that it is feasible to enhance performance with minimal effort.
Future Direction. We underscore several potential avenues for future research, detailed below.
  • Training data. Create a subset of training data with deceptive prompts similar to what we have in the MAD-Bench, create correct responses, and train the MLLM to resist deception.
  • Check consistency between image and prompt. Identify and interpret elements in the image, such as objects, colors, and spatial relationships. Then, analyze the question to understand its content and intent. Compare the two to identify any discrepancies before generating a response.
  • Focus on factual information. Ensure that the response sticks to information factually derived from the image. Refrain from making speculative assumptions or inferences that go beyond the scope of the image and the question.
6 Conclusion
In this study, we introduce MAD-Bench, a new benchmark comprising 850 image-prompt pairs, meticulously categorized into six distinct types of deceptive scenarios, to evaluate the robustness of state-of-the-art MLLMs against deceptive prompts. Our findings indicate a notable vulnerability in these models. Though GPT-4V achieves the best performance, it still exhibits substantial room for improvement. We hope our new benchmark can stimulate further research to enhance models’ resilience against deceptive prompts.
Limitation
When designing deceptive questions for our benchmark, we included a variety of categories to increase the diversity of the questions as a starting point. However, there are unlimited scenarios where MLLMs can be deceived. The additional piece of prompt added to boost model performance in Section 5 serves the purpose of demonstrating that simple efforts can improve the robustness of MLLMs in face of deceptive information. It is not optimized, thus not showing the maximum capability of this method.
Appendix A Appendix
Figure 9: Example of how MLLMs respond to deceptive prompts in the Count of Object category.
Figure 10: Example of how MLLMs respond to deceptive prompts in the Non-existent Object category.
Figure 11: Example of how MLLMs respond to deceptive prompts in the Object Attribute category.
Figure 11: Example of how MLLMs respond to deceptive prompts in the Object Attribute category.
Figure 12: Example of how MLLMs respond to deceptive prompts in the Scene Understanding category.
Figure 13: Example of how MLLMs respond to deceptive prompts in the Spatial Relationship category.
Figure 14: Example of how MLLMs respond to deceptive prompts in the Visual Confusion category.
Figure 15: Example of how MLLMs respond to deceptive prompts in the Visual Confusion category.
A.1 Examples of Responses from MLLMs to Deceptive Prompts
In Figures 9-15, we show examples of how MLLMs respond to deceptive prompts, and observe that there is a large gap between GPT-4V and other MLLMs on resisting deceptive prompts.
A.2 Prompts Used to Generate Deceptive Prompts using GPT-4
Count of Object
Illustrated in Figure 18.
Non-existent Object
Illustrated in Figure 18.
Object Attribute
Illustrated in Figure 18.
Scene Understanding
Illustrated in Figure 20.
Spatial Relationship
Illustrated in Figureble 20.
Figure 16: Prompt used to generate deceptive questions for the Count of Object category using GPT-4.
Figure 17: Prompt used to generate deceptive questions for the Non-existent Object category using GPT-4.
Figure 16: Prompt used to generate deceptive questions for the Count of Object category using GPT-4.
Figure 17: Prompt used to generate deceptive questions for the Non-existent Object category using GPT-4.
Figure 18: Prompt used to generate deceptive questions for the Object Attribute category using GPT-4.
Figure 19: Prompt used to generate deceptive questions for the Scene Understanding category using GPT-4.
Figure 19: Prompt used to generate deceptive questions for the Scene Understanding category using GPT-4.
Figure 20: Prompt used to generate deceptive questions for the Spatial Relationship category using GPT-4.
Visual Confusion
Due to the special nature of this category, all the prompts are human written instead of using GPT-4.
Figure 21: Prompts Used to Evaluate Responses from MLLM Using GPT-4.
A.3 Prompts Used to Evaluate Responses from MLLMs Using GPT-4
The prompts used to evaluate responses from the first five categories are listed in Figure 21. Due to the special nature of the Visual Confusion category, responses in this category are evaluated manually.